OpenStudy (anonymous):

F=9/5C=32 where F= 100.58 convert

OpenStudy (shiraz14):

Hi, I think you meant F = (9/5)*C + 32 --> C = (5/9)*(F - 32) Now, can you try converting F = 100.58 to C?

OpenStudy (anonymous):

i have no idea how to do this please help

OpenStudy (shiraz14):

Simply substitute F=100.58 into the equation above {C = (5/9)*(F - 32)} and you'll find the value of C when F=100.58. Can you do that?